OpenStudy (anonymous):

Can someone help me with finding zeros in a function? zeros of f(x)= 3x^2 - 3x - 6 / 3x-6

OpenStudy (anonymous):

\[\frac{ 3x^2-3x-6}{ 3x-6 }\]

terenzreignz (terenzreignz):

Wrong section. but try factoring the numerator.

OpenStudy (anonymous):

Factorize the numerator. \[\frac{ x ^{2} - 2x + x - 2 }{ x - 2 }\] \[\frac{(x - 2 ) ( x + 1) }{ x - 2 }\] Can you solve it now? :?